A point charge q is located at the center of a spherical shell of radius a that has a charge -q uniformly distributed on its surface.
Find the electric field for the following points:
(a) for all points outside the spherical shell
E = kq/r2
E = kq2/r2
E = q/4pr2
none of these
E = 0

(b) for a point inside the shell a distance r from the center
E = q/4pr2
E = kq2/r2
E = kq/r2
E = 0
none of these
Answer #1

A point charge q is located at the center of a spherical shell of radius a that has a charge -q uniformly distributed on its surface.


Find the electric field for the following points:
(a) for all points outside the spherical shell
E = kq/r2
E = kq2/r2
E = q/4pr2
none of these
E = 0

(b) for a point inside the shell a distance r from the center
E = q/4pr2
E = kq2/r2
E = kq/r2
E = 0
none of these
Answer #2
The enclosed charged in the spherical charge is zero due to induced charges.
hence elecricfield E = kq/r^2
                              = 0
now the for a point inside the shell
electric field E = kq/r^2